# PlainMamba: Improving Non-Hierarchical Mamba in Visual Recognition

### v1

@@ Abstract
We present PlainMamba: a simple non-hierarchical state space model (SSM) designed for general visual recognition. The recent Mamba model has shown how SSMs can be highly competitive with other architectures on sequential data and initial attempts have been made to apply it to images. In this paper, we further adapt the selective scanning process of Mamba to the visual domain, enhancing its ability to learn features from two-dimensional images by (i) a continuous 2D scanning process that improves spatial continuity by ensuring adjacency of tokens in the scanning sequence, and (ii) direction-aware updating which enables the model to discern the spatial relations of tokens by encoding directional information. Our architecture is designed to be easy to use and easy to scale, formed by stacking identical PlainMamba blocks, resulting in a model with constant width throughout all layers. The architecture is further simplified by removing the need for special tokens. We evaluate PlainMamba on a variety of visual recognition tasks including image classification, semantic segmentation, object detection, and instance segmentation. Our method achieves performance gains over previous non-hierarchical models and is competitive with hierarchical alternatives. For tasks requiring high-resolution inputs, in particular, PlainMamba requires much less computing while maintaining high performance. Code and models are available at https://github.com/ChenhongyiYang/PlainMamba
@@ 1 Introduction
Developing high-performing visual encoders has always been one of the most important goals in computer vision [39, 95, 73, 24, 80, 59, 23]. With high-quality visual features, a broad range of downstream tasks, such as semantic segmentation [107, 12, 94, 84], object recognition [80, 39, 95, 60] and detection [53, 38, 67] can be tackled with relative ease.
Early methods for extracting visual representations relied on hand-crafted features such as SIFT  [61] and SURF [4]. A big breakthrough then came with the adoption of convolutional neural networks (CNNs) that process images with local contexts and enforce spatial equivariance [47, 39, 73]. Recently, vision transformers (ViTs) [24] obviated the need for such enforced inductive biases in favour of learnable contexts that operate on image patches [82, 80, 59]. However, despite the overwhelming success of transformers and their self-attention mechanism [21, 8, 100], the quadratic cost of attention has proved to be an obstacle to further scaling such models.
@@
This has invigorated interest in state space models (SSMs) [90, 31, 34, 58, 108]. Due to their close ties to linear recurrent networks, SSMs have the benefits of potentially infinite context lengths while maintaining linear complexity in the input sequence length [31], offering substantial speedups compared to attention. However, it took several notable advances to make SSMs effective at learning competitive representations, including enforcing state space variables to be orthogonal basis projections [34]. The recent Mamba [31] architecture further aligned SSM-based models with modern transformers, such as making the state space variables input-dependent — much like queries, keys and values in self-attention. When being used for NLP, those designs led to a state space model that could scale to the sizes and performances of modern transformer-based LLMs [8, 81], while improving long-range context and inference speed.
@@
Figure 1: While hierarchical visual encoders may demonstrate superior accuracy on open-source visual recognition benchmarks, the plain non-hierarchical models have had more widespread use because of their simple structure. In this work, we investigate the potential of the plain Mamba model in visual recognition. Please note that the presence of lions in this figure is purely due to the first author’s predilection for cats, and is not in any way related to a former name (‘Simba’) of our model.
@@
There is now understandable interest in adapting the Mamba architecture to the visual domain [58, 49, 108]. However, before we start doing that, we need to think about under what guidelines should we design our new model. As we show in Figure 1, by examining the development of recently proposed visual encoders, we find that adding more inductive biases, e.g., hierarchical structure, to the plain model such as DeiT can indeed improve a model’s performance on open-source benchmarks like ImageNet. However, we should not ignore the fact that the plain ViT [24] is widely used by several popular vision foundation models [46, 56, 65, 63, 1], which suggests that simplicity in architecture design is key for multiple reasons. Firstly, maintaining a constant model width (i.e. non-hierarchical) makes it much easier to integrate features from multiple levels, as is common in dense prediction tasks such as semantic segmentation [46]. It also becomes easier to combine features across different modalities such as in CLIP [65] or LLaVa [56] or as parts of increasingly complex AI-powered systems. Furthermore, simpler components can be more easily optimized for hardware acceleration [17]. In addition, it has also been observed that the over-crafted models may lead to a significant gap between the performance on commonly used benchmarks and downstream tasks [26, 6]. This means benchmark performance may no longer reflect real-world usefulness, as over-engineering tends to increase model complexity and thus make it harder for others to re-use.
@@
Motivated by the above findings, we propose PlainMamba: a simple Mamba architecture for visual recognition. This model integrates ideas from CNNs, Transformers and novel SSM-based models with an aim to providing easy-to-use models for the vision modality. Compared to previous visual state space models [58, 108], we simplify the architecture by maintaining constant model width across all layers of the network via stacking identical blocks as well as removing the need for CLS tokens. This allows for easy scaling and model re-use, while achieving competitive performances.
Our contributions are as follows:
  • We propose a new visual state space model we call PlainMamba. This architecture improves and simplifies previous attempts at extending the Mamba architecture to the visual modality.
  • We improve the SSM block by adapting selective scanning to better process 2D spatial inputs, in two ways. (i) Our continuous 2D scanning approach ensures that the scanning sequence is spatially continuous to improve semantic continuity. (ii) Our direction-aware updating, inspired by positional encoding, allows the model to encode the directionality of each scanning order to further improve spatial context.
@@
  • We test our PlainMamba architecture using three different sizes (7M, 26M and 50M) and show how they perform competitively on a range of tasks, from ImageNet1K classification to semantic segmentation and object detection. Specifically, we show that PlainMamba outperforms its non-heretical counterparts, including SSMs and Transformers, while performing on par with the hierarchical competitors.
@@ 2.0.1 Visual Feature Extractors
How to effectively extract visual features from images has been a long-standing challenge in computer vision. In the early years of deep learning, CNNs [73, 47, 39, 95] dominated the model architecture landscape. Their induced spatial prior, through the use of convolutional filters, exploits the locality of visual features. Furthermore, stacking multiple layers increases their receptive field. Many different CNN backbone architectures have been proposed over the years [47, 44, 13], introducing new ways of exploiting spatial information [73, 95], building deeper models [39, 77], improving efficiency [71, 78, 66], adding multi-scale connections [69], scaling architectures [91], and introducing attention mechanisms [83, 5, 75, 85, 42, 10]. In recent years, ViTs have become a powerful tool for image modeling [24]. Compared to CNNs, they make fewer assumptions about data (feature locality [98], translation and scale invariance). By replacing the convolutional layers with self-attention modules, transformers can capture global relationships and have achieved state-of-the-art results on many common image benchmarks [20, 54, 107]. To adapt the original transformer architecture [82] for vision tasks, images are split into patches and converted into tokens before being fed into the transformer encoder. Within this framework, numerous works have focused on pushing the performance (e.g. LeViT, [30] combining transformer encoder layers and convolutions), or on reducing the costly quadratic complexity of self-attention [18, 16]. Another popular extension to ViT architectures has been the addition of hierarchical structures [86, 59, 97, 92, 27], similar to the multi-scale feature pyramids used in CNNs. The Swin Transformer [59], for instance, uses shifted windows to share feature information across scales. These multi-scale features are then used for a wide range of downstream tasks. Recent research has explored ways of using these hierarchical features within ViTs themselves [68, 36, 51, 23, 37, 9, 25, 48, 72]. Some works [59] have examined the use of multi-resolution features as attention keys and values to learn multi-scale information. However, these extensions add complexity to the model and make it harder to effectively use its features in later stages, thus hindering widespread adoption. Indeed, recent works [50, 100] return to the original ViT architecture, as its non-hierarchical nature greatly simplifies the use of its features. In particular, the plain ViT provides greater flexibility for pre-training and fine-tuning on different tasks.
2.0.2 State Space Models
@@
State Space Models (SSMs) have emerged as efficient alternatives to transformers and CNNs due to their ability to scale linearly with sequence length [33, 76]. SSMs transform the state space to effectively capture dependencies over extended sequences. To alleviate the initial computational cost of such models, S4 [32] enforced low-rank constraints on the state matrix and S5 [74] introduced parallel scanning to further improve efficiency. Furthermore, H3 [29] achieved competitive results on common benchmarks by improving the hardware utilization. Lastly, Mamba [31] parameterized the SSM matrices as functions of the input, thus allowing it to act as a learnable selection mechanism and providing greater flexibility. Follow-up works have extended selective SSMs for images [3, 70, 57, 89, 88] and videos [62] using a hierarchical structure [58] and bidirectional blocks [108], while Mamba-ND [49] introduces an architecture for multi-dimensional data. MambaIR [35] tackles image restoration, and Pan-Mamba [40] works on pan-sharpening. DiS [28] introduces SSMs to diffusion models by replacing the U-Net with an SSM backbone. While drawing inspiration from the above works, PlainMamba improves Mamba’s [31] selective SSM block by adding wider depth-wise convolutions. In contrast to the Cross-Scan Module (CSM) [58] and Mamba-ND [49], PlainMamba respects the spatio-sequential nature of image patches (see Figure 2). As opposed to [108], we do not use the CLS token.
2.0.3 Simplifying Visual Feature Extractors
@@
Simplifying and unifying existing methods is equally important as improving performance. Plain non-hierarchical architectures are robust, conceptually simpler, and scale better. In addition, they serve as a basis for future research. ViTs [24] remove the pyramid structure of CNNs by converting images into patched tokens. This way, they easily adapt the transformer architecture for visual tasks. Another trick that stems form sequence modeling is the usage of CLS tokens for prediction, which have proven to be unnecessary for visual tasks [103]. FlexiVit [7] unified into a single architecture images with different input resolutions, and GPViT [100] improved feature resolution with a non-hierarchical transformer. Similarly, ConvNext [60] introduced a simple CNN model that competed with state-of-the-art transformer methods. Other works, like MLP-Mixer [79] and follow-up works [41], have introduced simple architectures using only multi-layer perceptrons. The plain non-hierarchical ViT [24] has served as a simple building block for many diverse tasks. SAM [46] uses a pre-trained ViT as image encoder with minimal changes for image segmentation at large scale. DinoV2 [63, 19] uses a ViT to learn general-purpose visual features by pretraining models on curated datasets with self-supervision. Similarly, the image encoder for the CLIP [65] model consists of a basic ViT with minor modifications, allowing image-text representations to be learned with a contrastive objective. DALLE-2 [1] incorporates a ViT image encoder to extract visual features that are used for text-conditional image generation. LlaVA [56, 55] combines a vision encoder (pretrained ViT from CLIP) and an LLM for vision-language tasks. Inspired by the versatility and flexibility of ViT, this work aims to continue in the direction of simplifying feature extractors while maintaining strong performance by introducing PlainMamba: a Simple Mamba architecture.

### v2

@@ Abstract
We present PlainMamba: a simple non-hierarchical state space model (SSM) designed for general visual recognition. The recent Mamba model has shown how SSMs can be highly competitive with other architectures on sequential data and initial attempts have been made to apply it to images. In this paper, we further adapt the selective scanning process of Mamba to the visual domain, enhancing its ability to learn features from two-dimensional images by (i) a continuous 2D scanning process that improves spatial continuity by ensuring adjacency of tokens in the scanning sequence, and (ii) direction-aware updating which enables the model to discern the spatial relations of tokens by encoding directional information. Our architecture is designed to be easy to use and easy to scale, formed by stacking identical PlainMamba blocks, resulting in a model with constant width throughout all layers. The architecture is further simplified by removing the need for special tokens. We evaluate PlainMamba on a variety of visual recognition tasks, achieving performance gains over previous non-hierarchical models and is competitive with hierarchical alternatives. For tasks requiring high-resolution inputs, in particular, PlainMamba requires much less computing while maintaining high performance. Code and models are available at: https://github.com/ChenhongyiYang/PlainMamba.
@@ 1 Introduction
Developing high-performing visual encoders has always been one of the most important goals in computer vision [He et al.(2016)He, Zhang, Ren, and Sun, Xie et al.(2016)Xie, Girshick, Dollár, Tu, and He, Simonyan and Zisserman(2015), Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby, Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou, Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo, Dong et al.(2022)Dong, Bao, Chen, Zhang, Yu, Yuan, Chen, and Guo]. With high-quality visual features, a broad range of downstream tasks, such as semantic segmentation [Zhou et al.(2017)Zhou, Zhao, Puig, Fidler, Barriuso, and Torralba, Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Xiao et al.(2018)Xiao, Liu, Zhou, Jiang, and Sun, Wang et al.(2023)Wang, Kim, Feris, and Betke], object recognition [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou, He et al.(2016)He, Zhang, Ren, and Sun, Xie et al.(2016)Xie, Girshick, Dollár, Tu, and He, Liu et al.(2022)Liu, Mao, Wu, Feichtenhofer, Darrell, and Xie] and detection [Lin et al.(2017b)Lin, Goyal, Girshick, He, and Dollar, He et al.(2017)He, Gkioxari, Dollar, and Girshick, Ren et al.(2015)Ren, He, Girshick, and Sun] can be tackled with relative ease. Early methods for extracting visual representations relied on hand-crafted features such as SIFT  [Lowe(2004)] and SURF [Bay et al.(2006)Bay, Tuytelaars, and Van Gool]. A big breakthrough then came with the adoption of convolutional neural networks (CNNs) that process images with local contexts and enforce spatial equivariance [Krizhevsky et al.(2012)Krizhevsky, Sutskever, and Hinton, He et al.(2016)He, Zhang, Ren, and Sun, Simonyan and Zisserman(2015)]. Recently, vision transformers (ViTs) [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] obviated the need for such enforced inductive biases in favour of learnable contexts that operate on image patches [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin, Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou, Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo]. However, despite the overwhelming success of transformers and their self-attention mechanism [Devlin et al.(2019)Devlin, Chang, Lee, and Toutanova, Brown et al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal, Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan, Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess, Clark, Berner, McCandlish, Radford, Sutskever, and Amodei, Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang], the quadratic cost of attention has proved to be an obstacle to further scaling such models.
@@
This has invigorated interest in state space models (SSMs) [Williams et al.(2007)Williams, Lawrence, et al., Gu and Dao(2023), Gu et al.(2023)Gu, Johnson, Timalsina, Rudra, and Ré, Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang]. Due to their close ties to linear recurrent networks, SSMs have the benefits of potentially infinite context lengths while maintaining linear complexity in the input sequence length [Gu and Dao(2023)], offering substantial speedups compared to attention. However, it took several notable advances to make SSMs effective at learning competitive representations, including enforcing state space variables to be orthogonal basis projections [Gu et al.(2023)Gu, Johnson, Timalsina, Rudra, and Ré]. The recent Mamba [Gu and Dao(2023)] architecture further aligned SSM-based models with modern transformers, such as making the state space variables input-dependent — much like queries, keys and values in self-attention. When being used for NLP, those designs led to a state space model that could scale to the sizes and performances of modern transformer-based LLMs [Brown et al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal, Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan, Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess, Clark, Berner, McCandlish, Radford, Sutskever, and Amodei, Touvron et al.(2023)Touvron, Lavril, Izacard, Martinet, Lachaux, Lacroix, Rozière, Goyal, Hambro, Azhar, et al.], while improving inference effiency.
@@
Figure 1: While hierarchical visual encoders may demonstrate superior accuracy on open-source visual recognition benchmarks, the plain non-hierarchical models have had more widespread use because of their simple structure. We investigate the potential of the plain Mamba model in visual recognition.
@@
There is now understandable interest in adapting the Mamba architecture to the visual domain [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Li et al.(2024)Li, Singh, and Grover, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang]. However, before we start doing that, we need to think about under what guidelines should we design our new model. As we show in Figure 1, by examining the development of recently proposed visual encoders, we find that adding more inductive biases, e.g., hierarchical structure, to the plain model such as DeiT can indeed improve a model’s performance on open-source benchmarks like ImageNet. However, we should not ignore the fact that the plain ViT [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] is widely used by several popular vision foundation models [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick, Liu et al.(2023b)Liu, Li, Wu, and Lee, Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, Howes, Huang, Xu, Sharma, Li, Galuba, Rabbat, Assran, Ballas, Synnaeve, Misra, Jegou, Mairal, Labatut, Joulin, and Bojanowski, et al(2022)], which suggests that simplicity in architecture design is key for multiple reasons. Firstly, maintaining a constant model width (i.e. non-hierarchical) makes it much easier to integrate features from multiple levels, as is common in dense prediction tasks such as semantic segmentation [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick]. It also becomes easier to combine features across different modalities such as in CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] or LLaVa [Liu et al.(2023b)Liu, Li, Wu, and Lee] or as parts of increasingly complex AI-powered systems. Furthermore, simpler components can be more easily optimized for hardware acceleration [Dao et al.(2022a)Dao, Fu, Ermon, Rudra, and Ré]. In addition, it has also been observed that the over-crafted models may lead to a significant gap between the performance on commonly used benchmarks and downstream tasks [Ericsson et al.(2021)Ericsson, Gouk, and Hospedales, Beyer et al.(2020)Beyer, Hénaff, Kolesnikov, Zhai, and Oord]. This means benchmark performance may no longer reflect real-world usefulness, as over-engineering tends to increase model complexity and thus make it harder for others to re-use.
@@
Motivated by the above findings, we propose PlainMamba: a simple Mamba architecture for visual recognition. This model integrates ideas from CNNs, Transformers and novel SSM-based models with an aim to providing easy-to-use models for the vision modality. Compared to previous visual state space models [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang], we simplify the architecture by maintaining constant model width across all layers of the network via stacking identical blocks as well as removing the need for CLS tokens. This allows for easy scaling and model re-use, while achieving competitive performances.
@@
Our contributions are as follows: (1) We propose a new visual state space model we call PlainMamba. This architecture improves and simplifies previous attempts at extending the Mamba architecture to the visual modality. (3) We improve the SSM block by adapting selective scanning to better process 2D spatial inputs, in two ways. (i) Our continuous 2D scanning approach ensures that the scanning sequence is spatially continuous to improve semantic continuity. (ii) Our direction-aware updating, inspired by positional encoding, allows the model to encode the directionality of each scanning order to further improve spatial context. (3)We test our PlainMamba architecture using three different sizes (7M, 26M and 50M) and show how they perform competitively on a range of tasks, from ImageNet1K classification to semantic segmentation and object detection. Specifically, we show that PlainMamba outperforms its non-heretical counterparts, including SSMs and Transformers, while performing on par with the hierarchical competitors.
@@ 2 Related Work
Visual Feature Extractors. How to effectively extract visual features from images has been a long-standing challenge in computer vision. In the early years of deep learning, CNNs [Simonyan and Zisserman(2015), Krizhevsky et al.(2012)Krizhevsky, Sutskever, and Hinton, He et al.(2016)He, Zhang, Ren, and Sun, Xie et al.(2016)Xie, Girshick, Dollár, Tu, and He] dominated the model architecture landscape. Their induced spatial prior, through the use of convolutional filters, exploits the locality of visual features. Furthermore, stacking multiple layers increases their receptive field. Many different CNN backbone architectures have been proposed over the years [Krizhevsky et al.(2012)Krizhevsky, Sutskever, and Hinton, Huang et al.(2017)Huang, Liu, van der Maaten, and Weinberger, Chollet(2016)], introducing new ways of exploiting spatial information [Simonyan and Zisserman(2015), Xie et al.(2016)Xie, Girshick, Dollár, Tu, and He], building deeper models [He et al.(2016)He, Zhang, Ren, and Sun, Szegedy et al.(2015)Szegedy, Liu, Jia, Sermanet, Reed, Anguelov, Erhan, Vanhoucke, and Rabinovich], improving efficiency [Sandler et al.(2018)Sandler, Howard, Zhu, Zhmoginov, and Chen, Tan and Le(2019), Radosavovic et al.(2020)Radosavovic, Kosaraju, Girshick, He, and Dollár], adding multi-scale connections [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox], scaling architectures [Woo et al.(2023)Woo, Debnath, Hu, Chen, Liu, Kweon, and Xie], and introducing attention mechanisms [Wang et al.(2017)Wang, Jiang, Qian, Yang, Li, Zhang, Wang, and Tang, Bello(2021), Srinivas et al.(2021)Srinivas, Lin, Parmar, Shlens, Abbeel, and Vaswani, Wang et al.(2020)Wang, Wu, Zhu, Li, Zuo, and Hu, Hu et al.(2018)Hu, Shen, and Sun, Chen et al.(2019)Chen, Fan, Xu, Yan, Kalantidis, Rohrbach, Yan, and Feng]. In recent years, ViTs have become a powerful tool for image modeling [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby]. Compared to CNNs, they make fewer assumptions about data (feature locality [Xu et al.(2021b)Xu, Zhang, Zhang, and Tao], translation and scale invariance). By replacing the convolutional layers with self-attention modules, transformers can capture global relationships and have achieved state-of-the-art results on many common image benchmarks [Deng et al.(2009)Deng, Dong, Socher, Li, Li, and Fei-Fei, Lin et al.(2014)Lin, Maire, Belongie, Hays, Perona, Ramanan, Dollár, and Zitnick, Zhou et al.(2017)Zhou, Zhao, Puig, Fidler, Barriuso, and Torralba]. To adapt the original transformer architecture [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin] for vision tasks, images are split into patches and converted into tokens before being fed into the transformer encoder. Within this framework, numerous works have focused on pushing the performance (e.g. LeViT, [Graham et al.(2021)Graham, El-Nouby, Touvron, Stock, Joulin, Jégou, and Douze] combining transformer encoder layers and convolutions), or on reducing the costly quadratic complexity of self-attention [Dao et al.(2022b)Dao, Fu, Ermon, Rudra, and Ré, Dao(2024)]. Another popular extension to ViT architectures has been the addition of hierarchical structures [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao, Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo, Xu et al.(2021a)Xu, Xu, Chang, and Tu, Wu et al.(2021)Wu, Xiao, Codella, Liu, Dai, Yuan, and Zhang, Fan et al.(2021)Fan, Xiong, Mangalam, Li, Yan, Malik, and Feichtenhofer], similar to the multi-scale feature pyramids used in CNNs. The Swin Transformer [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo], for instance, uses shifted windows to share feature information across scales. These multi-scale features are then used for a wide range of downstream tasks. Recent research has explored ways of using these hierarchical features within ViTs themselves [Ren et al.(2022)Ren, Zhou, He, Feng, and Wang, Guo et al.(2022)Guo, Han, Wu, Tang, Chen, Wang, and Xu, Li et al.(2022b)Li, Wu, Fan, Mangalam, Xiong, Malik, and Feichtenhofer, Dong et al.(2022)Dong, Bao, Chen, Zhang, Yu, Yuan, Chen, and Guo, Hatamizadeh et al.(2022)Hatamizadeh, Yin, Kautz, and Molchanov, Chen et al.(2022a)Chen, Panda, and Fan, d’Ascoli et al.(2021)d’Ascoli, Touvron, Leavitt, Morcos, Biroli, and Sagun, Lee et al.(2022)Lee, Kim, Willette, and Hwang, Shi(2024)]. Some works [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] have examined the use of multi-resolution features as attention keys and values to learn multi-scale information. However, these extensions add complexity to the model and make it harder to effectively use its features in later stages, thus hindering widespread adoption. Indeed, recent works [Li et al.(2022a)Li, Mao, Girshick, and He, Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang] return to the original ViT architecture, as its non-hierarchical nature greatly simplifies the use of its features. In particular, the plain ViT provides greater flexibility for pre-training and fine-tuning on different tasks.
Figure 2: (a) The overall architecture of the proposed PlainMamba. PlainMamba does not have a hierarchical structure, it instead stacks $L$ identical PlainMamba block to form the main network. For image classification, it uses global average pooling instead of the CLS to gather global information. (b) Architecture of PlainMamba block, which is similar to the Mamba [Gu and Dao(2023)] block where the selective scanning is combined with a gated MLP. (c) The proposed Direction-Aware Updating, where a series of learnable parameters $\mathbf{\Theta}_{k}$ are combined with the data-dependent updating parameters to explicitly inject relative 2D positional information into the selective scanning process.
@@
State Space Models. State Space Models (SSMs) have emerged as efficient alternatives to transformers and CNNs due to their ability to scale linearly with sequence length [Gu et al.(2021)Gu, Johnson, Goel, Saab, Dao, Rudra, and Re, Sun et al.(2023)Sun, Dong, Huang, Ma, Xia, Xue, Wang, and Wei]. SSMs transform the state space to effectively capture dependencies over extended sequences. To alleviate the initial computational cost of such models, S4 [Gu et al.(2022)Gu, Goel, and Re] enforced low-rank constraints on the state matrix and S5 [Smith et al.(2023)Smith, Warrington, and Linderman] introduced parallel scanning to further improve efficiency. Furthermore, H3 [Fu et al.(2023)Fu, Dao, Saab, Thomas, Rudra, and Ré] achieved competitive results on common benchmarks by improving the hardware utilization. Lastly, Mamba [Gu and Dao(2023)] parameterized the SSM matrices as functions of the input, thus allowing it to act as a learnable selection mechanism and providing greater flexibility. Follow-up works have extended selective SSMs for images [Baron et al.(2023)Baron, Zimerman, and Wolf, Ruan and Xiang(2024), Liu et al.(2024a)Liu, Yang, Zhou, Xi, Yu, Yu, Liang, Shi, Zhang, Zheng, et al., Wang et al.(2024)Wang, Zheng, Zhang, Cui, and Li, Wang and Ma(2024), Ma et al.(2024)Ma, Li, and Wang, Pei et al.(2024)Pei, Huang, and Xu, Liu et al.(2024b)Liu, Yang, Zhou, Xi, Yu, Yu, Liang, Shi, Zhang, Zheng, et al.] and videos [Nguyen et al.(2022)Nguyen, Goel, Gu, Downs, Shah, Dao, Baccus, and Ré] using a hierarchical structure [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] and bidirectional blocks [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang], while Mamba-ND [Li et al.(2024)Li, Singh, and Grover] introduces an architecture for multi-dimensional data. MambaIR [Guo et al.(2024)Guo, Li, Dai, Ouyang, Ren, and Xia] tackles image restoration, and Pan-Mamba [He et al.(2024)He, Cao, Yan, Li, Xie, Zhang, and Zhou] works on pan-sharpening. DiS [Fei et al.(2024)Fei, Fan, Yu, and Huang] introduces SSMs to diffusion models by replacing the U-Net with an SSM backbone. While drawing inspiration from the above works, PlainMamba improves Mamba’s [Gu and Dao(2023)] selective SSM block by adding wider depth-wise convolutions. In contrast to the Cross-Scan Module (CSM) [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] and Mamba-ND [Li et al.(2024)Li, Singh, and Grover], PlainMamba respects the spatio-sequential nature of image patches (see Figure 2). As opposed to [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang], we do not use the CLS token.
@@
Simplifying Visual Feature Extractors. Simplifying and unifying existing methods is equally important as improving performance. Plain architectures are robust, conceptually simpler, and scale better. ViTs [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] remove the pyramid structure of CNNs by converting images into patched tokens. This way, they easily adapt the transformer architecture for visual tasks. Another trick that stems form sequence modeling is the usage of CLS tokens for prediction, which have proven to be unnecessary for visual tasks [Zhai et al.(2022)Zhai, Kolesnikov, Houlsby, and Beyer]. FlexiVit [Beyer et al.(2023)Beyer, Izmailov, Kolesnikov, Caron, Kornblith, Zhai, Minderer, Tschannen, Alabdulmohsin, and Pavetic] unified into a single architecture images with different input resolutions, and GPViT [Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang] improved feature resolution with a non-hierarchical transformer. Similarly, ConvNext [Liu et al.(2022)Liu, Mao, Wu, Feichtenhofer, Darrell, and Xie] introduced a simple CNN model that competed with state-of-the-art transformer methods. Other works, like MLP-Mixer [Tolstikhin et al.(2021)Tolstikhin, Houlsby, Kolesnikov, Beyer, Zhai, Unterthiner, Yung, Steiner, Keysers, Uszkoreit, et al.] and follow-up works [Hou et al.(2022)Hou, Jiang, Yuan, Cheng, Yan, and Feng], have introduced simple architectures using only multi-layer perceptrons. The plain non-hierarchical ViT [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] has served as a simple building block for many diverse tasks. SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick] uses a pre-trained ViT as image encoder with minimal changes for image segmentation at large scale. DinoV2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, Howes, Huang, Xu, Sharma, Li, Galuba, Rabbat, Assran, Ballas, Synnaeve, Misra, Jegou, Mairal, Labatut, Joulin, and Bojanowski, Darcet et al.(2023)Darcet, Oquab, Mairal, and Bojanowski] uses a ViT to learn general-purpose visual features by pretraining models on curated datasets with self-supervision. Similarly, the image encoder for the CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] model consists of a basic ViT with minor modifications, allowing image-text representations to be learned with a contrastive objective. DALLE-2 [et al(2022)] incorporates a ViT image encoder to extract visual features that are used for text-conditional image generation. LlaVA [Liu et al.(2023b)Liu, Li, Wu, and Lee, Liu et al.(2023a)Liu, Li, Li, and Lee] combines a vision encoder (pretrained ViT from CLIP) and an LLM for vision-language tasks.